OpenStudy (anonymous):

Calculate |-5 + 12i|.

OpenStudy (anonymous):

@SolomonZelman @iambatman @Jhannybean

hartnn (hartnn):

\(\Large |a+bi| = \sqrt{a^2+b^2}\)

hartnn (hartnn):

here a= -5 b=12

OpenStudy (dtan5457):

\[\sqrt{(-5)^2+(12)^2}=?\]

OpenStudy (anonymous):

so root 25+144?

hartnn (hartnn):

yes go on

OpenStudy (anonymous):

13?

hartnn (hartnn):

\(\sqrt{25+144}\)

OpenStudy (dtan5457):

yes

hartnn (hartnn):

13 is correct :)

OpenStudy (anonymous):

thank you! :)

hartnn (hartnn):

welcome ^_^